OpenStudy (anonymous):

cos2x= Square root of 2 minus cos2x. Solving for all solutions between [0, 2pie).

OpenStudy (shamim):

is it\[\cos2x=\sqrt{2-\cos2x}\]

OpenStudy (anonymous):

No

OpenStudy (anonymous):

\[\cos 2x= \sqrt{2}- \cos2x\]

OpenStudy (shamim):

ok

OpenStudy (shamim):

or,\[2\cos2x=\sqrt{2}\]

OpenStudy (shamim):

or,\[\cos 2 x=\frac{ 1 }{ \sqrt{2} }\]

OpenStudy (shamim):

or,\[\cos 2x=\cos \frac{ \pi }{ 4 }\]

OpenStudy (shamim):

i think u will b able do the rest